In a certain store, there is a .02 probability that the scanned price in the bar code scanner will not match the advertised price. The cashier scans 844 items.

(a-1) What is the expected number of mismatches? (Round your answer to the nearest whole number.)
  
Expected number            

(a-2) What is the standard deviation? (Use your rounded number for the expected number of mismatches for the calculation of standard deviation. Round your final answer to 4 decimal places.)
  
Standard deviation            

(b) What is the approximate normal probability of at least 10 mismatches? (Round the z-value to 2 decimal places. Use Appendix C-2 to find probabilities. Round your final answer to 4 decimal places.)
  
Probability            

(c) What is the approximate normal probability of more than 22 mismatches? (Round the z-value to 2 decimal places. Use Appendix C-2 to find probabilities. Round your final answer to 4 decimal places.)
  
Probability            
Answer #1


Given that, n=844, p=0.02 a-1) The expected number of mismatches is, A =np =844*0.02 =16.88 =17 a-2) The standard deviation i= 0.9671 The approximate normal probability of more than 22 mismatches is, P(XBinomial > 22) = P(X.Normal > 22+0.5) = P( Vera